Write an equation in standard form using only integers for the line described.
A line parallel to 5x-7y=35 (6,1)
(the correct answer is 5x-7y=23)

Answer by richwmiller(17219)

5x-7y=23 is only one of many possible answers.
5x-7y=35
-7y=-5x+35
y=-5x/-7-5
y=5x/7-5
slope is 5/7
y=mx+b
1=5/7*6+b
1=30/7+b
1-30/7=b
-23/7=b
y=5x/7=23/7
We must use integers not fractions
7y=5x-23
23=5x-7y
5x-7y=23